OpenStudy (anonymous):

is this true or false? {1} \(\in\) {1}

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

1 belongs to the list containing 1 ? yeah it does... or did you mean 1 \(\in\) I ? where I is an Integer set?

OpenStudy (anonymous):

If you cant read the latex i dont know why it showed up before \(\in\) = E

OpenStudy (freckles):

The element in question is {1} is the element {1} contained in the set containing 1?

OpenStudy (anonymous):

no

OpenStudy (freckles):

then you have your answer what if it was \[\left\{ 1 \right\} \in \left\{ \left\{ 1 \right\} \right\}\] true or false

OpenStudy (anonymous):

true, funny thing is is i have that problem too

OpenStudy (freckles):

lol that is funny

hartnn (hartnn):

freckles is the author of the problems you're practicing :P

OpenStudy (anonymous):

wow this class im in i already marked {1} \(\in\) {1} as false i forgot i was checking my answer that quick, on this one though {{1}} \(\subsetq\) {1, {1}} is false

OpenStudy (anonymous):

i have no idea why that did not work... subsetq

OpenStudy (anonymous):

...

OpenStudy (anonymous):

ok now its working again

hartnn (hartnn):

right, thats false

OpenStudy (anonymous):

ok i think i got the rest right then thanks